# Fedor V. Fomin

### v1

@@ Abstract
The connection between Hamiltonicity and the independence numbers of graphs has been a fundamental aspect of Graph Theory since the seminal works of the 1960s. This paper presents a novel algorithmic perspective on these classical problems. Our contributions are twofold.
First, we establish that a wide array of problems in undirected graphs, encompassing problems such as Hamiltonian Path and Cycle, Path Cover, Largest Linkage, and Topological Minor Containment are fixed-parameter tractable ($\operatorname{{\sf FPT}}$) parameterized by the independence number of a graph. To the best of our knowledge, these results mark the first instances of $\operatorname{{\sf FPT}}$ problems for such parameterization.
Second, we extend the algorithmic scope of the Gallai-Milgram theorem. The original theorem by Gallai and Milgram, asserts that for a graph $G$ with the independence number $\alpha(G)$, the vertex set of $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths. We show that determining whether a graph can be covered by fewer than $\alpha(G)-k$ vertex-disjoint paths is $\operatorname{{\sf FPT}}$ parameterized by $k$.
@@ 1 Introduction
A significant portion of parameterized algorithms deals with optimization problems on sparse graphs when specific graph parameters characterize the sparsity and structure of the graph. Such parameters include treewidth, treedepth, or minimum vertex cover size. However, the parameterized complexity of many problems needs to be understood more when it comes to the parameters describing the graph’s density, such as its independence number, that is, the size of the maximum independent set.
The initial step in exploring algorithmic properties of graphs with bounded independence numbers was taken by Ovetsky Fradkin and Seymour in their work [FS15]. They demonstrated that a generalization of the edge-disjoint path problem, which they call “fusion”, can be solved in polynomial time on digraphs with a constant independence number. The work of Ovetsky Fradkin and Seymour, as well as the follow-up research [${}^{+}$20, MSSZ23], indicates that the class of digraphs with bounded independence numbers exhibits sufficient structure for algorithmic exploitation.
In this paper, we leverage the structure of undirected graphs with bounded independence numbers to develop parameterized algorithms. Our approach is applicable to a wide range of combinatorial problems, including Hamiltonicity and path covers in undirected graphs. The interplay between Hamiltonicity and independence numbers is a fundamental theme in Graph Theory, stemming from the classic theorem of Gallai–Milgram [GM60]. Surprisingly, the computational complexity of the Graph Hamiltonicity problem remained unresolved, even for graphs with an independence number of $\alpha(G)=3$, until the recent work by Jedlicková and Kratochvíl [JK23]. Jedlicková and Kratochvíl have demonstrated that not only Graph Hamiltonicity but also more general $-\ell$-Linkage problems (which involve connecting $\ell$ pairs of vertices with vertex-disjoint paths covering all the graph’s vertices) can be solved in polynomial time when the independence number $\alpha(G)$ of a graph is constant.
The algorithms proposed by Jedlicková and Kratochvíl for Graph Hamiltonicity is an XP algorithm for parameter $\alpha(G)$. This means that on an $n$-vertex graph $G$, the running times of these algorithms are the form of $n^{f(\alpha(G))}$ for some function $f$. To the best of our knowledge, no fundamental NP-hard problem has been found to be $\operatorname{{\sf FPT}}$ parameterized by $\alpha(G)$. (The only trivial exception is coloring a graph $G$ in a constant number of colors $q$ because every yes-instance has at most $q\cdot\alpha(G)$ vertices.) This leads Jedlicková and Kratochvíl in  [JK23] to an intriguing question: What is the parameterized complexity of the Hamiltonian Path problem (and many other related problems) parameterized by $\alpha(G)$?
@@
We find the situation concerning parameterization by $\alpha(G)$ unusual: Most algorithmic findings in structural parameterization require that computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$. For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ parameterized by the treewidth of a graph due to the following reasoning. First, for an input integer $k$, there are (either exact or approximation) algorithms that either correctly decide that the treewidth of the input graph is more than $k$ or construct a tree decomposition of width $\mathcal{O}(k)$ in time $f(k)\cdot n^{\mathcal{O}(1)}$ [Bod96, Kor21, RS86]. Then, the tree decomposition is used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$. In contrast, determining whether $\alpha(G)\geq k$ is W[1]-hard with parameter $k$ [DF99]. Moreover, there is compelling evidence that the problem does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22]. To overcome the intractability of the independence number, we design algorithms with the following properties. For a given parameter $k$, the algorithm either solves an optimization problem at hand (Hamiltonicity, Path Cover, etc.) in $\operatorname{{\sf FPT}}$ time with respect to $k$, or it correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$. Such algorithms are what Raghavan and Spinrad referred to as robust [RS03].
Our contribution. To state formally our first main result, we need the following definitions.
@@
Let $H$, $G$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)\subseteq V(G)$, such that all vertices of $M\setminus T$ have degree $2$. We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to a graph obtained from $M$ by dissolving in $M$ all vertices of $M\setminus T$. Here, dissolving a vertex of degree two means deleting it while making its neighbors adjacent. We further say that for the injective mapping $f:V(H)\to V(G)$ induced by the isomorphism above, $(M,f)$ is a topological minor embedding (TM-embedding) of $H$ in $G$. The size of a TM-model $(M,T)$ and of a TM-embedding $(M,f)$ is the number of vertices in $M$.
@@
We also consider a more general variant of TM-embedding. In this variant, every vertex of $H$ is assigned a list of vertices of $G$ where it could be mapped.
@@
Our first result is the theorem establishing that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$ parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$.
@@
List TM-embedding encompasses several vital problems in graph algorithms and Theorem 1 implies several corollaries.
@@
  • When graph $H$ consists of a path $P_{2}$ on two vertices whose lists are all vertices of $V(G)$, then this is the problem of finding a longest path in $G$. If this path contains $|V(G)|$ vertices, this is a Hamiltonian path. For $H=C_{3}$, we get the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$. By Theorem 1, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$. This positively answers the open question from  [JK23].
@@
  • In Path Cover, the task is to cover all vertices of $G$ by at most $p$ vertex-disjoint paths. By Gallai–Milgram’s theorem [GM60], if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$. We can solve the problem by taking $H$ to be disjoint unions of at most $p$ paths $P_{1}$ and $P_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$. Thus, Theorem 1 yields an algorithm of running time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ solving this problem. For Cycle Cover, that is when the graph $H$ consists of at most $p$ disjoint copies of $C_{3}$, we have the running time of $f(\alpha(G)+p)\cdot|G|^{\mathcal{O}(1)}$.
@@
  • Jedlicková and Kratochvíl in  [JK23] defined the following generalization of Path Cover. In the $\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$. The task is to connect these pairs by disjoint paths that altogether traverse all vertices of the graph. For $\ell=1$ this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path. $\ell$-Linkage could be encoded as List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$. Let the vertices of these $P_{2}$’s be $\{x_{i},y_{i}\}$, $i\in\{1,\dots,\ell\}$. Then the list mapping $L$ assigns $\{s_{i}\}$ to $x_{i}$ and $\{t_{i}\}$ to $y_{i}$. By Theorem 1, $\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$, resolving another open question of Jedlicková and Kratochvíl. By playing with different graphs $H$ in the List TM-embedding and Theorem 1, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.

### v2

@@ Abstract
The classic theorem of Gallai and Milgram (1960) asserts that for every graph $G$, the vertex set of $G$ can be partitioned into at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the maximum size of an independent set in $G$. The proof of Gallai–Milgram’s theorem is constructive and yields a polynomial-time algorithm that computes a covering of $G$ by at most $\alpha(G)$ vertex-disjoint paths.
We prove the following algorithmic extension of Gallai–Milgram’s theorem for undirected graphs: determining whether an undirected graph can be covered by fewer than $\alpha(G)-k$ vertex-disjoint paths is fixed-parameter tractable ($\mathsf{FPT}$) when parameterized by $k$. More precisely, we provide an algorithm that, for an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, runs in time
$2^{k^{O(k^{4})}}\cdot n^{O(1)},$
and outputs a path cover $\mathcal{P}$ of $G$. Furthermore, it:
  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover,
  • or outputs, together with $\mathcal{P}$, an independent set of size $\lvert\mathcal{P}\rvert+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
A key subroutine in our algorithm is an $\mathsf{FPT}$ algorithm, parameterized by $\alpha(G)$, for deciding whether $G$ contains a Hamiltonian path. This result is of independent interest—prior to our work, no polynomial-time algorithm for deciding Hamiltonicity was known even for graphs with independence number at most 3. Moreover, the techniques we develop apply to a wide array of problems in undirected graphs, including Hamiltonian Cycle, Path Cover, Largest Linkage, and Topological Minor Containment. We show that all these problems are $\mathsf{FPT}$ when parameterized by the independence number of the graph.
@@ 1 Introduction
The Gallai–Milgram Theorem, published by Tibor Gallai and Arthur Milgram in 1960 [GM60], is a cornerstone of structural graph theory and can be viewed as a far-reaching generalization of several classical results, such as Dilworth’s Theorem [Dil50], concerning the covering of graphs by simpler substructures. In its original form, the theorem states that every (directed or undirected) graph $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the size of a maximum independent set in $G$.
Beyond its immediate implications, the Gallai–Milgram Theorem belongs to a broader line of research in graph theory focused on connections between Hamiltonicity and independence. This theme can be traced back to the seminal work of Rédei [Réd34], who showed that every tournament (a directed graph with $\alpha(G)=1$) contains a directed Hamiltonian path. Further investigations into the relationships among stability number, connectivity, and Hamiltonicity began with Nash-Williams [NW71] and Chvátal–Erdős [CE72], and have continued in later works (see, e.g., [AFG83, BBVR89, Jac87, Koi94, BAH88, BT07, HJ90, Tho01, FS09]).
In this paper, we present a novel algorithmic perspective on this classic theme in graph theory: the interplay among the Gallai–Milgram theorem, Hamiltonicity, and the independence number.
1.1 Our contribution.
Known constructive proofs of the Gallai–Milgram theorem can be used to construct in polynomial time a covering of the vertex set of a graph $G$ by at most $\alpha(G)$ vertex-disjoint paths. Our following theorem presents a significant generalization of this algorithmic result for undirected graphs: for an integer parameter $k$, deciding whether a graph $G$ contains a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf FPT}}$, parameterized by $k$.
Theorem 1.
There is an algorithm that, given an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, in time $2^{k^{\mathcal{O}(k^{4})}}\cdot n^{\mathcal{O}(1)}$ outputs a cover of $V(G)$ by a family $\mathcal{P}$ of vertex-disjoint paths and, furthermore,
Theorem 1 aligns with a broader trend in parameterized algorithms that seeks algorithmic analogues of classic theorems in extremal graph theory. Examples include algorithmic extensions of Poljak–Turzík’s theorem [MPSS14], Dirac’s theorem [FGSS22a], and Hajnal–Szemerédi’s theorem by Gan, Han, and Hu [GHH23]. As with Theorem 1, each of these results provides an $\operatorname{{\sf FPT}}$ algorithm parameterized “above” or “beyond” a known combinatorial bound. A distinguishing feature of Theorem 1, however, is that in previous work the relevant combinatorial bounds (e.g., the maximum vertex degree) were efficiently computable, whereas determining whether $\alpha(G)\geq k$ is $\operatorname{{\sf W}}[1]$-hard with respect to $k$ [DF99] and likely does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22]. Consequently, we cannot simply require $\alpha(G)$ to be part of the input, which makes the existence of an $\operatorname{{\sf FPT}}$ algorithm parameterized below $\alpha(G)$ all the more surprising.
As a special case of Theorem 1, let $k=\alpha(G)-1$. In this scenario, the problem reduces to determining whether all vertices of $G$ can be covered by a single path—i.e., whether $G$ has a Hamiltonian path. Consequently, Theorem 1 implies that deciding whether $G$ contains a Hamiltonian path is $\operatorname{{\sf FPT}}$ when parameterized by $\alpha(G)$. The construction of such an $\operatorname{{\sf FPT}}$ algorithm for Hamiltonicity is a key component of the proof of Theorem 1. Moreover, the techniques developed to achieve this result extend well beyond Hamiltonicity and apply to a broader class of problems.
@@
Similar to Theorem 1, the situation concerning parameterization by $\alpha(G)$ is unusual: Most algorithmic findings in structural parameterization require that computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$. For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ parameterized by the treewidth of a graph due to the following reasoning. First, for an input integer $k$, there are (either exact or approximation) algorithms that either correctly decide that the treewidth of the input graph is more than $k$ or construct a tree decomposition of width $\mathcal{O}(k)$ in time $f(k)\cdot n^{\mathcal{O}(1)}$ [Bod96, Kor21, RS86]. Then, the tree decomposition is used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$. In contrast, this time we deal with the parameter that does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22].
Our second theorem provides an $\operatorname{{\sf FPT}}$ algorithm for computing topological minors in a graph, parameterized by its independence number.
Let us recall that a graph $H$ is called a topological minor of a graph $G$ if there exists a subgraph of $G$ that is isomorphic to a subdivision of $H$. A subdivision of a graph is formed by replacing one or more edges with simple paths (i.e., inserting additional vertices along those edges). Thus, saying that $H$ is a topological minor of $G$ means that $G$ contains a subgraph $M$ that represents a “subdivided” version of $H$, which we refer to as a topological minor model $M$ of $H$ in $G$. Many interesting problems can be expressed in terms of topological minors. For example, the problem of computing a longest cycle in a graph is equivalent to computing the largest model $M$ of a graph $H=C_{3}$, the cycle on three vertices. More formally,
@@
Let $H$, $G$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)$ be a vertex set such that all vertices of $M\setminus T$ have degree $2$. We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to a graph obtained from $M$ by dissolving in $M$ all vertices of $M\setminus T$. Here, dissolving a vertex of degree two means deleting it while making its neighbors adjacent. We further say that for the injective mapping $f:V(H)\to V(G)$ induced by the isomorphism above, $(M,f)$ is a topological minor embedding (TM-embedding) of $H$ in $G$. The size of a TM-model $(M,T)$ and of a TM-embedding $(M,f)$ is the number of vertices in $M$.
@@
To capture topological minor models containing specific terminal vertices, like finding a longest cycle containing a specified terminal vertex set $T$, we also consider a more general variant of TM-embedding. In this variant, every vertex of $H$ is assigned a list of vertices of $G$ where it could be mapped.
@@
The following theorem establishes that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$ parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$. The algorithm, for a given parameter $k$, either computes a TM-embedding of maximum size in $\operatorname{{\sf FPT}}$ time with respect to $k$, or correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$. Such algorithms are what Raghavan and Spinrad referred to as robust [RS03]. More precisely,
@@
To our knowledge, neither list TM-embeddings nor the Maximum List TM-Embedding problem have been studied so far. However, this problem is a natural generalization of several vital problems in graph algorithms. In particular, Theorem 2 implies the following corollaries.
@@
  • When graph $H$ consists of a path $P_{2}$ on two vertices whose lists are all vertices of $V(G)$, then this is the problem of finding a longest path in $G$. If this path contains $|V(G)|$ vertices, this is a Hamiltonian path. For $H=C_{3}$, we have the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$. By Theorem 2, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$.
@@
  • In Path Cover, the task is to cover all vertices of $G$ by at most $p$ vertex-disjoint paths. By Gallai–Milgram’s theorem, if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$. We can solve the problem by taking $H$ to be disjoint unions of at most $p$ paths $P_{1}$ and $P_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$. Thus, Theorem 2 yields an algorithm of running time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ solving this problem. Of course, it also follows from Theorem 1. For Cycle Cover, that is when the graph $H$ consists of at most $p$ disjoint copies of $C_{3}$, we have the running time $f(\alpha(G)+p)\cdot|G|^{\mathcal{O}(1)}$.
@@
  • Consider the following generalization of Path Cover. In the $-\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$. The task is to connect these pairs by disjoint paths that altogether traverse all vertices of the graph. For $\ell=1$ this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path. $-\ell$-Linkage could be encoded as List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$. Let the vertices of these $P_{2}$’s be $\{x_{i},y_{i}\}$, $i\in\{1,\dots,\ell\}$. Then the list mapping $L$ assigns $\{s_{i}\}$ to $x_{i}$ and $\{t_{i}\}$ to $y_{i}$. By Theorem 2, $-\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$. By playing with different graphs $H$ in the List TM-embedding and Theorem 2, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.